OpenStudy (anonymous):

the Width of a rectangle is 5 inches greater that 3 times its length. If the area of the rectangle is 22 square inches, what is the width of the rectangle in inches?

OpenStudy (anonymous):

3L+5+3L+5+L+L=22 8L+10=22 8L=12 L=12/8 L=1.5 square inches w=3L+5 w=3(1.5)+5 w=9.5 square inches w=width l=length

OpenStudy (anonymous):

okay??

OpenStudy (campbell_st):

length = l width = 3l+5 area of the rectangle A =l(3l+5) or \[22= 3l^2 + 5l\] the quadratic is \[3l^2 + 5l - 22 = 0\]' solve the quadratic( it can be factroised) (l -2)(3l + 11) l = 2, -11 only consider the positive as you are finding a measurement. the length of the rectangle is 2inches the width is 11 inches

OpenStudy (precal):